A 0.20 kg baseball is traveling at 40 m/s toward the batter. The ball is hit by the bat with a force of 200N, and is now headed toward the pitcher at 25 m/s. How long was the bat in contact with the ball? 


1
Expert's answer
2021-02-19T10:30:03-0500
"F\Delta t=m(v_f-v_i),""\Delta t=\dfrac{F}{m|(v_f-v_i)|},""\Delta t=\dfrac{200\ N}{0.20\ kg\cdot|(-25\ \dfrac{m}{s}-40\ \dfrac{m}{s})|}=15.4\ s."